# Dissipative currents and transport coefficients in relativistic spin hydrodynamics

### v1

@@ I Introduction
Relativistic spin hydrodynamics has emerged as a possible theory of phenomenological interest for the description of the strongly interacting Quark Gluon Plasma in relativistic heavy ion collisions, driven by experimental evidence of spin polarization of the produced particles  [1, 2]; for a recent review see, for example, Ref. [3]. For this reason, there has been a lot of interest in literature on this topic  [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37], and yet its physical meaning is still under debate. In this theory, pseudo-gauge invariance of a combined transformation of energy-momentum and spin tensor is broken [38], and, unlike in the most usual formulation of relativistic hydrodynamics, the description of the fluid dynamics may need to include a non-vanishing spin tensor if, for instance, the system features a separation of scales in the relaxation times of spin and momentum degrees of freedom [10]. However, the relation between the spin tensor and the quantum field operators is not unique; in other words, the pseudo-gauge fixing remains arbitrary.
@@
A key problem in relativistic spin hydrodynamics is the derivation of dissipative currents as functions of gradients of thermo-hydrodynamic fields (four-velocity, temperature, chemical potentials, spin potential or suitable combinations thereof). Ambiguities may arise between different formulations depending on how these dissipative currents are decomposed, which can significantly impact the final physical results. In this work, we take advantage of a recent study [39] to provide general forms of dissipative currents up to first order in gradients of thermo-hydrodynamic fields. Our method is quite general and provides the maximal number of transport coefficients subject to the so-called matching condition of vanishing projections of the dissipative part of the currents onto the four-velocity field and the condition of positive entropy production.
@@ II Dissipative currents and their decomposition
In the quantum-statistical description of a relativistic fluid, the entropy production is obtained by calculating the divergence of the out-of-equilibrium entropy current, which is constructed from the local equilibrium density operator $\widehat{\rho}_{\rm LE}$, and the von Neumann entropy $S=-\mathrm{Tr}\big{[}\widehat{\rho}_{\rm LE}\log\widehat{\rho}_{\rm LE}\big{]}$. The resulting expression for the entropy production rate reads [39],
@@
Here, $\xi_{\mu\nu}=\partial_{(\mu}\beta_{\nu)}$ is the thermal shear tensor, $\varpi_{\mu\nu}=\partial_{[\nu}\beta_{\mu]}$ is the thermal vorticity tensor, and the fields $\beta_{\mu},~{}\zeta,~{}\Omega_{\mu\nu}$ are the thermal velocity four-vector $u_{\mu}/T$, the chemical potential over temperature $\mu/T$, and the spin potential over temperature $\omega_{\mu\nu}/T$, respectively.
@@
the choice of $n^{\mu}$ is arbitrary and corresponds to the freedom of choosing a frame in the classical formulation of relativistic hydrodynamics. A possible choice for $n^{\mu}$ is the fluid 4-velocity $u^{\mu}$ (as shown in Fig. 1) motivated by the fact that with this choice one recovers matching conditions similar to those in Landau’s or Eckart’s formulations of dissipative hydrodynamics without a spin tensor. However, in the quantum relativistic statistical formulation which led to Eq. (4), the condition (6) applied to the four-velocity field implies a restriction which is absent in the classical formulation and should be kept in mind when comparing the final results. Particularly, if the velocity field has a finite vorticity (like at global thermodynamic equilibrium with non-vanishing angular velocity), $n$ cannot coincide with $u$ and should be defined otherwise (see the discussion in Ref. [40]).
@@
and we will use them to express the dissipative currents at first order in the gradient expansion. It should be noted that in other studies the spin potential $\Omega$ (and the thermal vorticity as well) is itself considered as ”first-order quantity” in the gradients so that $\partial\Omega$ is neglected as being a ”second order quantity”. We do not make such an assumption, thus we tacitly assume $\Omega$ and $\varpi$ to be not necessarily small. On the other hand, we will neglect the anisotropy induced by a potentially large thermal vorticity and spin potential when decomposing the tensor-valued coefficients, as will be made clear shortly.
@@
where the traceless symmetric projector $\Delta_{\mu\nu,\alpha\beta}\equiv(\Delta_{\mu\alpha}\Delta_{\nu\beta}+\Delta_{\nu\alpha}\Delta_{\mu\beta})/2-\Delta_{\mu\nu}\Delta_{\alpha\beta}/3$ was used. Substituting the decomposed tensor coefficients $H,K,L,M,\dots,W$ back into the dissipative currents in Eqs. (8)-(11), and subsequently into the entropy production rate (4), results in an explicit expression of $\partial_{\mu}s^{\mu}$, which will be analyzed further in the following.
@@ III Matching conditions and semi-positivity constraints
After taking these 34 constraints into account, we are left with $98-34=64$ undetermined coefficients.
@@
where we have introduced the comoving and spatial derivatives $D\equiv u_{\mu}\partial^{\mu}$ and $\nabla^{\mu}\equiv\Delta^{\mu\nu}\partial_{\nu}$, the expansion scalar $\theta\equiv\partial_{\mu}u^{\mu}$, the heat flow $\mathcal{J}_{h}^{\mu}\equiv Du^{\mu}-(1/T)\nabla^{\mu}T$, and the shear tensor $\sigma^{\mu\nu}\equiv\Delta^{\mu\nu}_{\alpha\beta}\partial^{\alpha}u^{\beta}$. The following gradient term, namely the difference between the spin potential and the thermal vorticity, $\Omega_{\mu\nu}-\varpi_{\mu\nu}$, is antisymmetric and can be decomposed as,
@@
where the electric-like part is $\mathcal{E}^{\mu}\equiv Tu_{\nu}(\Omega^{\nu\mu}-\varpi^{\nu\mu})$, and the magnetic-like part reads $\mathcal{B}^{\mu}\equiv T\epsilon^{\alpha\beta\gamma\mu}u_{\gamma}(\Omega_{\alpha\beta}-\varpi_{\alpha\beta})$. The decomposition of the next gradient contribution, i.e., the derivative of the of the chemical potential over temperature, is given by
@@
Finally, the derivative of the spin potential is decomposed as follows, with details of the derivation relegated to Appendix A,
@@
[…]
@@
We note that $\partial_{\mu}s^{\mu}$ is a Lorentz scalar which is invariant under reflections as well as spatial rotations in the rest frame associated to $u^{\mu}$. Regarding the thermo-hydrodynamic gradient terms appearing in Eqs. (III)-(26), we may classify them according to their behavior under this class of transformations, as listed in Table 6. In order to ensure all of the terms appearing in the entropy production rate are (parity-even) scalars, some of the coefficients in Eqs. (III)-(26) have to be parity-odd while others are parity-even. For simplicity, we set all parity-odd coefficients to zero. This further eliminates 28 of the coefficients, bringing down the total number to $64-28=36$. Then, we can use the information in Table 6 to divide the entropy production rate $\partial_{\mu}s^{\mu}$ into the following six terms,
@@
These constraints include 13 equalities, which reduce the number of independent coefficients to $36-13=23$. The remaining quantities are further constrained by the inequalities introduced in Eqs. (48)–(53), which can be classified as follows: 11 inequalities demand that a certain coefficient has a definite sign, 6 inequalities set a bound for the product of two coefficients, and 1 inequality constrains the product of three coefficients. Taking all of these into account, we can construct the final form of the dissipative currents, as will be done in the following.
@@ IV Final expressions and comparison with previous results
and the relevant transport coefficients are summarized in Table 8. The coefficients $\kappa_{h}$ and $\gamma_{\phi}$ have been previously obtained in Ref. [7] and named boost heat conductivity and rotational viscosity, denoted as $\lambda$ and $\gamma$ respectively. It should be noted that the current associated with the coefficient $\kappa_{h}$ does not precisely match that of $\lambda$ in Ref. [7] due to the choice of the matching condition. Nevertheless, they basically have the same physical meaning. Studies on the rotational viscosity coefficient have been also conducted in Refs. [41, 42]. We thus identify three new transport coefficients: $\kappa_{\zeta}$, $\kappa_{\mathcal{G}},$ and $\gamma_{\Xi}$, which appear due to the inclusion of the gradient of the spin potential and the finite chemical potential.
@@
with transport coefficients listed in Table 10. The coefficients $\chi_{Z},\chi_{F},\chi_{\Xi}$ were identified in Refs. [18, 29, 35]; most coefficients related to the derivatives of the spin potential were obtained in Refs. [8, 36]. In addition to the constraints in Tables 7-10, we have another requirement:
@@
where $\mu\equiv\zeta T$ is the chemical potential, $\rho$ the proper energy density, $p$ the pressure, $n$ the charge density, $S^{\mu\nu}\equiv u_{\lambda}{\cal S}^{\lambda,\mu\nu}$ is the “spin density” and $\omega_{\mu\nu}=T\Omega_{\mu\nu}$. Nevertheless, the differential relation in eq. (59) may not be fulfilled, as it was pointed out in Ref. [39], so that using the equation (4) as a starting point appears to be a safer and more general method.
@@ V Application to quantum kinetic theory
Since the second law of thermodynamics is a macroscopic statement, any (first-order) microscopic theory has to comply with it. One such example is the formulation of spin hydrodynamics from quantum kinetic theory [26, 43, 44, 45]. Before being able to properly apply the results derived in the preceding sections, we have to clarify the power-counting used in quantum kinetic theory, which is referred to as the $\hbar$-gradient expansion and denotes an expansion around the classical limit. In particular, in the standard counting that treats spin-related quantities as small quantum corrections, one has $\hbar\partial_{\lambda}S^{\lambda\mu\nu}\sim T_{A}^{\mu\nu}\sim\mathcal{O}(\hbar^{2})$. On the other hand, $j^{\mu}\sim T_{S}^{\mu\nu}\sim\mathcal{O}(1)$ are of zeroth order, with corrections emerging at second order. These corrections have not been computed in Refs. [26, 43, 44, 45], so we are a bit limited in our analysis. We also have to revise the conditions given in Eqs. (14) and (15). The reason is that, since in the aforementioned references the currents $\delta j^{\mu}$ and $\delta T_{S}^{\mu\nu}$ are only evaluated up to first order in $\hbar$, the matching conditions are $u_{\mu}\delta j^{(0)\mu}=0$ and $u_{\mu}\delta T_{S}^{(0)\mu\nu}=0$. Here, quantities marked with an index $(0)$ denote the contributions of zeroth order in quantum corrections. The form of the currents is then
@@
For the comparison, we focus on the formulation derived in Ref. [45]. Therein, the information on dissipation is included in the bulk viscous pressure $\Pi$, the diffusion current $n^{\mu}$, the shear-stress tensor $\pi^{\mu\nu}$, and a traceless rank-two tensor $\mathfrak{t}^{\mu\nu}$ appearing in the spin tensor. These quantities are dynamical and follow relaxation-type equations [detailed in Eq. (2) of Ref. [45]]. To find their Navier-Stokes values, i.e., relate them algebraically to the first- order quantities at our disposal, we only keep those terms in the respective equations of motion that are linear in the first-order tensors. [^2] In the context of this work, we also treat the components of $\partial^{\lambda}\Omega^{\mu\nu}$ as being of first order when finding the Navier-Stokes limit of $\mathfrak{t}^{\mu\nu}$. Also, we identified $\mathcal{F}_{S}^{\mu\nu}\simeq\nabla^{\langle\mu}\kappa^{\nu\rangle}$. In particular, keeping to the notation of Ref. [45] and denoting potentially ambiguous quantities with a subscript “kin”, we have
@@
with $\delta j^{(2)\mu}_{\text{kin}}$ and $\delta T_{S,\text{kin}}^{(2)\mu\nu}$ denoting the second-order contributions that have not been computed in Ref. [45]. Furthermore, $\sigma$ is the spin of the particle, and $I_{nq}$, $J_{nq}$ are thermodynamic integrals, which for Boltzmann statistics read
@@
We emphasize that the fact that many coefficients vanish is a consequence of the simple truncation employed in Sec. IX of Ref. [45]. When using a higher-order truncation, a lot of the coefficients will become nonzero, allowing for further nontrivial inequalities to be checked. These investigations are left for future work. Considering the subset of coefficients that are nonzero, the inequalities (V) become
@@
The first three inequalities are well-known from the case of standard hydrodynamics, while the following three appear only for nonzero spin. Note that the conditions $\Gamma^{(\kappa)}\geq 0$ and $\Gamma^{(\omega)}\geq 0$ can be related to the damping of the spin waves, i.e., the rate at which the spin potential approaches the thermal vorticity [44].
@@
$10^{-2}$$10^{-1}$$10^{0}$$10^{1}$$10^{-8}$$10^{-6}$$10^{-4}$$10^{-2}$$10^{0}$$z$$\eta_{\text{kin}}[1/(TG^{2})]$$\kappa_{\text{kin}}[1/(T^{2}G^{2})]$$\zeta_{\text{kin}}[1/(TG^{2})]$$10^{-2}$$10^{-1}$$10^{0}$$10^{1}$$10^{-8}$$10^{-6}$$10^{-4}$$10^{-2}$$10^{0}$$z$$z^{2}\Gamma^{(\kappa)}[G^{2}T^{8}]$$z^{2}\Gamma^{(\omega)}[G^{2}T^{8}]$$\ell_{\mathfrak{t}\kappa}[1/(TG^{2})]$

### v2

@@ I Introduction
Relativistic spin hydrodynamics has emerged as a possible theory of phenomenological interest for the description of the strongly interacting Quark Gluon Plasma in relativistic heavy ion collisions, driven by experimental evidence of spin polarization of the produced particles  [STAR:2017ckg, STAR:2019erd]; for a recent review see, for example, Ref. [Becattini:2024uha]. For this reason, there has been a lot of interest in literature on this topic  [Becattini:2011zz, Florkowski:2017ruc, Hattori:2019ahi, Hattori:2019lfp, Gallegos:2020otk, Li:2020eon, Hongo:2021ona, Fukushima:2020ucl, Gallegos:2021bzp, Montenegro:2017rbu, Montenegro:2018bcf, Bhadury:2020puc, Garbiso:2020puw, Weickgenannt:2020aaf, Shi:2020htn, She:2021lhe, Cartwright:2021qpp, Bhadury:2021oat, Gallegos:2022jow, Peng:2021ago, Liu:2021uhn, Fu:2021pok, Daher:2022xon, Weickgenannt:2022zxs, Hu:2022azy, Daher:2022wzf, Biswas:2023qsw, Daher:2024ixz, Daher:2024bah, Kiamari:2023fbe, Weickgenannt:2023bss, Kumar:2023ojl, Ren:2024pur, Yang:2024duc, Drogosz:2024gzv, Drogosz:2024lkx, Bhadury:2025boe, Florkowski:2024bfw, Dey:2024crk, Dey:2024cwo, Fang:2025aig, Buzzegoli:2024mra, Buzzegoli:2025zud, Tiwari:2024trl], and yet its physical meaning is still under debate. In this theory, pseudo-gauge invariance of a combined transformation of energy-momentum and spin tensor is broken [Becattini:2018duy], and, unlike in the most usual formulation of relativistic hydrodynamics, the description of the fluid dynamics may need to include a non-vanishing spin tensor if, for instance, the system features a separation of scales in the relaxation times of spin and momentum degrees of freedom [Hongo:2021ona]. However, the relation between the spin tensor and the quantum field operators is not unique; in other words, the pseudo-gauge fixing remains arbitrary.
@@
A key problem in relativistic spin hydrodynamics is the derivation of dissipative currents as functions of gradients of thermo-hydrodynamic fields (four-velocity, temperature, chemical potentials, spin potential or suitable combinations thereof). Ambiguities may arise between different formulations depending on how these dissipative currents are decomposed, which can significantly impact the final physical results. In this work, we take advantage of a recent study [Becattini:2023ouz] to provide general forms of dissipative currents up to first order in gradients of thermo-hydrodynamic fields. Our method is quite general and provides the maximal number of transport coefficients subject to the so-called matching condition of vanishing projections of the dissipative part of the currents onto the four-velocity field and the condition of positive entropy production.
@@ II Dissipative currents and their decomposition
In the quantum-statistical description of a relativistic fluid, the entropy production is obtained by calculating the divergence of the out-of-equilibrium entropy current, which is constructed from the local equilibrium density operator $\widehat{\rho}_{\rm LE}$, and the von Neumann entropy $S=-\mathrm{Tr}\big[\widehat{\rho}_{\rm LE}\log\widehat{\rho}_{\rm LE}\big]$. The resulting expression for the entropy production rate reads [Becattini:2023ouz],
@@
Here, $\xi_{\mu\nu}=\partial_{(\mu}\beta_{\nu)}$ is the thermal shear tensor, $\varpi_{\mu\nu}=\partial_{[\nu}\beta_{\mu]}$ is the thermal vorticity tensor, and the fields $\beta_{\mu},~\zeta,~\Omega_{\mu\nu}$ denote, respectively, the thermal velocity four-vector $u_{\mu}/T$, the reduced chemical potential, defined as the chemical potential $\mu$ over temperature $T$, and the reduced spin potential, defined as the spin potential $\omega_{\mu\nu}$ over temperature $T$.
@@
the choice of $n^{\mu}$ is arbitrary and corresponds to the freedom of choosing a frame in the classical formulation of relativistic hydrodynamics. A possible choice for $n^{\mu}$ is the fluid 4-velocity $u^{\mu}$ (as shown in Fig. 1) motivated by the fact that with this choice one recovers matching conditions similar to those in Landau’s or Eckart’s formulations of dissipative hydrodynamics without a spin tensor. However, in the quantum relativistic statistical formulation which led to Eq. (4), the condition (6) applied to the four-velocity field implies a restriction which is absent in the classical formulation and should be kept in mind when comparing the final results. Particularly, if the velocity field has a finite vorticity (like at global thermodynamic equilibrium with non-vanishing angular velocity), $n$ cannot coincide with $u$ and should be defined otherwise (see the discussion in Ref. [Becattini:2014yxa]).
@@
and we will use them to express the dissipative currents at first order in the gradient expansion. The gradient counting order of the term $(\Omega_{\mu\nu}-\omega_{\mu\nu})$ is not assumed but rather imposed by the entropy production, ensuring that all terms in Eq. (4) are of the same gradient order. Consequently, we treat both $\Omega$ and $\varpi$ as quantities that are not small. However, it should be noted that in other studies the reduced spin potential $\Omega$ (and the thermal vorticity as well) is itself considered as ”first-order quantity” in the gradients so that $\partial\Omega$ is neglected as being a ”second order quantity”. We do not make such an assumption, thus we tacitly assume $\Omega$ and $\varpi$ to be not necessarily small. On the other hand, we will neglect the anisotropy induced by a potentially large thermal vorticity and spin potential when decomposing the tensor-valued coefficients, as will be made clear shortly.
@@
where the traceless symmetric projector $\Delta_{\mu\nu,\alpha\beta}\equiv(\Delta_{\mu\alpha}\Delta_{\nu\beta}+\Delta_{\nu\alpha}\Delta_{\mu\beta})/2-\Delta_{\mu\nu}\Delta_{\alpha\beta}/3$ was used. The tensor coefficients $H^{\mu\nu\rho\sigma}$ (together with $K^{\mu\nu\rho},\dots,W^{\mu\lambda\nu\rho\sigma\tau}$) are decomposed in equilibrium, which is isotropic in the local co-moving frame of the fluid. In this case, the decompositions do not contain terms involving $\Omega$ because such terms breaks the local rotational invariance. Substituting the decomposed tensor coefficients $H,K,L,M,\dots,W$ back into the dissipative currents in Eqs. (8)-(11), and subsequently into the entropy production rate (4), results in an explicit expression of $\partial_{\mu}s^{\mu}$, which will be analyzed further in the following.
@@ III Matching conditions and semi-positivity constraints
After taking these 34 constraints into account, we are left with 9834=6498-34=64 undetermined coefficients.
@@
where we have introduced the comoving and spatial derivatives $D\equiv u_{\mu}\partial^{\mu}$ and $\nabla^{\mu}\equiv\Delta^{\mu\nu}\partial_{\nu}$, the expansion scalar $\theta\equiv\partial_{\mu}u^{\mu}$, the heat flow $\mathcal{J}_{h}^{\mu}\equiv Du^{\mu}-(1/T)\nabla^{\mu}T$, and the shear tensor $\sigma^{\mu\nu}\equiv\Delta^{\mu\nu}_{\alpha\beta}\partial^{\alpha}u^{\beta}$. The following gradient term, namely the difference between the reduced spin potential and the thermal vorticity, $\Omega_{\mu\nu}-\varpi_{\mu\nu}$, is antisymmetric and can be decomposed as,
@@
where the electric-like part is $\mathcal{E}^{\mu}\equiv Tu_{\nu}(\Omega^{\nu\mu}-\varpi^{\nu\mu})$, and the magnetic-like part reads $\mathcal{B}^{\mu}\equiv T\epsilon^{\alpha\beta\gamma\mu}u_{\gamma}(\Omega_{\alpha\beta}-\varpi_{\alpha\beta})$. The decomposition of the next gradient contribution, i.e., the derivative of the reduced chemical potential over temperature, is given by
@@
Finally, the derivative of the reduced spin potential is decomposed as follows, with details of the derivation relegated to Appendix A,
@@
$\theta$ $\mathcal{H}$ $\mathcal{J}_{h}^{\mu}$ $\mathcal{B}^{\mu}$ $\sigma^{\mu\nu}$ $\mathcal{I}^{\mu\nu}_{S}$
@@
We note that $\partial_{\mu}s^{\mu}$ is a Lorentz scalar which is invariant under reflections as well as spatial rotations in the rest frame associated to $u^{\mu}$. Regarding the thermo-hydrodynamic gradient terms appearing in Eqs. (III)-(26), we may classify them according to their behavior under this class of transformations, as listed in Table 6. In order to ensure all of the terms appearing in the entropy production rate are (parity-even) scalars, some of the coefficients in Eqs. (III)-(26) have to be parity-odd while others are parity-even. For simplicity, we set all parity-odd coefficients to zero. This further eliminates 28 of the coefficients, bringing down the total number to 6428=3664-28=36. Then, we can use the information in Table 6 to divide the entropy production rate $\partial_{\mu}s^{\mu}$ into the following six terms,
@@
These constraints include 13 equalities, which reduce the number of independent coefficients to 3613=2336-13=23. The remaining quantities are further constrained by the inequalities introduced in Eqs. (48)–(53), which can be classified as follows: 11 inequalities demand that a certain coefficient has a definite sign, 6 inequalities set a bound for the product of two coefficients, and 1 inequality constrains the product of three coefficients. Taking all of these into account, we can construct the final form of the dissipative currents, as will be done in the following.
@@ IV Final expressions and comparison with previous results
and the relevant transport coefficients are summarized in Table 8. The coefficients $\kappa_{h}$ and $\gamma_{\phi}$ have been previously obtained in Ref. [Hattori:2019lfp] and named boost heat conductivity and rotational viscosity, denoted as $\lambda$ and $\gamma$ respectively. It should be noted that the current associated with the coefficient $\kappa_{h}$ does not precisely match that of $\lambda$ in Ref. [Hattori:2019lfp] due to the choice of the matching condition. Nevertheless, they basically have the same physical meaning. Studies on the rotational viscosity coefficient have been also conducted in Refs. [Hongo:2022izs, Hidaka:2023oze]. We thus identify three new transport coefficients: $\kappa_{\zeta}$, $\kappa_{\mathcal{G}},$ and $\gamma_{\Xi}$, which appear due to the inclusion of the gradient of the spin potential and the finite chemical potential.
@@
with transport coefficients listed in Table 10. The coefficients $\chi_{Z},\chi_{F},\chi_{\Xi}$ were identified in Refs. [She:2021lhe, Biswas:2023qsw, Drogosz:2024gzv]; most coefficients related to the derivatives of the spin potential were obtained in Refs. [Gallegos:2020otk, Dey:2024cwo]. In addition to the constraints in Tables 7-10, we have another requirement:
@@
where $\mu\equiv\zeta T$ is the chemical potential, $\rho$ the proper energy density, $p$ the pressure, $n$ the charge density, $S^{\mu\nu}\equiv u_{\lambda}{\cal S}^{\lambda,\mu\nu}$ is the “spin density” and $\omega_{\mu\nu}=T\Omega_{\mu\nu}$. Nevertheless, the differential relation in eq. (59) may not be fulfilled, as it was pointed out in Ref. [Becattini:2023ouz], so that using the equation (4) as a starting point appears to be a safer and more general method.
@@ V Application to quantum kinetic theory
Since the second law of thermodynamics is a macroscopic statement, any (first-order) microscopic theory has to comply with it. One such example is the formulation of spin hydrodynamics from quantum kinetic theory [Weickgenannt:2022zxs, Weickgenannt:2022qvh, Wagner:2024fhf, Wagner:2024fry]. Before being able to properly apply the results derived in the preceding sections, we have to clarify the power-counting used in quantum kinetic theory, which is referred to as the $\hbar$-gradient expansion and denotes an expansion around the classical limit. In particular, in the standard counting that treats spin-related quantities as small quantum corrections, one has $\hbar\partial_{\lambda}S^{\lambda\mu\nu}\sim T_{A}^{\mu\nu}\sim\mathcal{O}(\hbar^{2})$. On the other hand, $j^{\mu}\sim T_{S}^{\mu\nu}\sim\mathcal{O}(1)$ are of zeroth order, with corrections emerging at second order. These corrections have not been computed in Refs. [Weickgenannt:2022zxs, Weickgenannt:2022qvh, Wagner:2024fhf, Wagner:2024fry], so we are a bit limited in our analysis. We also have to revise the conditions given in Eqs. (14) and (15). The reason is that, since in the aforementioned references the currents $\delta j^{\mu}$ and $\delta T_{S}^{\mu\nu}$ are only evaluated up to first order in $\hbar$, the matching conditions are $u_{\mu}\delta j^{(0)\mu}=0$ and $u_{\mu}\delta T_{S}^{(0)\mu\nu}=0$. Here, quantities marked with an index $(0)$ denote the contributions of zeroth order in quantum corrections. The form of the currents is then
@@
For the comparison, we focus on the formulation derived in Ref. [Wagner:2024fry]. Therein, the information on dissipation is included in the bulk viscous pressure $\Pi$, the diffusion current $n^{\mu}$, the shear-stress tensor $\pi^{\mu\nu}$, and a traceless rank-two tensor $\mathfrak{t}^{\mu\nu}$ appearing in the spin tensor. These quantities are dynamical and follow relaxation-type equations [detailed in Eq. (2) of Ref. [Wagner:2024fry]]. To find their Navier-Stokes values, i.e., relate them algebraically to the first- order quantities at our disposal, we only keep those terms in the respective equations of motion that are linear in the first-order tensors. [^2] In the context of this work, we also treat the components of $\partial^{\lambda}\Omega^{\mu\nu}$ as being of first order when finding the Navier-Stokes limit of $\mathfrak{t}^{\mu\nu}$. Also, we identified $\mathcal{F}_{S}^{\mu\nu}\simeq\nabla^{\langle\mu}\kappa^{\nu\rangle}$. In particular, keeping to the notation of Ref. [Wagner:2024fry] and denoting potentially ambiguous quantities with a subscript “kin”, we have
@@
with $\delta j^{(2)\mu}_{\text{kin}}$ and $\delta T_{S,\text{kin}}^{(2)\mu\nu}$ denoting the second-order contributions that have not been computed in Ref. [Wagner:2024fry]. Furthermore, $\sigma$ is the spin of the particle, and $I_{nq}$, $J_{nq}$ are thermodynamic integrals, which for Boltzmann statistics read
@@
We emphasize that the fact that many coefficients vanish is a consequence of the simple truncation employed in Sec. IX of Ref. [Wagner:2024fry]. When using a higher-order truncation, a lot of the coefficients will become nonzero, allowing for further nontrivial inequalities to be checked. These investigations are left for future work. Considering the subset of coefficients that are nonzero, the inequalities (V) become
@@
The first three inequalities are well-known from the case of standard hydrodynamics, while the following three appear only for nonzero spin. Note that the conditions $\Gamma^{(\kappa)}\geq 0$ and $\Gamma^{(\omega)}\geq 0$ can be related to the damping of the spin waves, i.e., the rate at which the reduced spin potential approaches the thermal vorticity [Wagner:2024fhf].
@@
[…]